# Liquid He Detectors for Neutrons

### v1

@@ Abstract
The He(n,p) process is excellent for neutron detection between thermal and 4 MeV because of the high cross section and near-complete energy transfer from the neutron to the proton. Traditional gaseous He detectors using this process typically have high levels of radiogenic backgrounds so that they cannot measure the small neutron fluxes present in underground labs for dark matter experiments. I propose a cryogenic liquid He detector that can be designed with tiny radiogenic backgrounds and efficiently measure neutron fluxes in low-flux environments.
@@
Low background underground environments, like SNOLAB, are home to many dark matter searches due to their minimization of non-signal backgrounds. The underground environments typically mitigate neutrons from cosmic rays (cosmogenic) very well, but there remains a neutron flux in the environment, primarily due to radiogenic sources–spontaneous fission or (,n) from the surrounding material [1]. The levels of these backgrounds are so low that they are difficult to measure, but still pose a threat to dark matter detectors as a background. My work aims to develop a prototype liquid He detector to measure the neutron backgrounds in low flux environments. The high cross section reaction He(n,p)H in combination with the high He number density in the liquid state will allow these low fluxes to be measured by this prototype [2].
@@
Figure 1 shows the neutron flux as a function of energy at SNOLAB. It was estimated from a combination of studies done in the past. First was a Monte Carlo approach targeting radiogenic neutrons from the cavern walls [3], this constrained the shape of the distribution for 6–10 MeV neutrons. The overall normalization for the “fast” neutrons was taken from the SNOLAB handbook [4]–the value was 4000 n/m/day. The thermal flux–measured for SNO–is taken to come from a Maxwell-Boltzman spectrum with overall normalization 4144.949.8105.3 n/m/day [5]. The flux in the area between these two measurements is generally unmeasured and has been interpolated with a simple power-law function in Fig. 1, . Our prototype detector will measure the flux in this energy region which is especially important for low-mass dark matter searches.
@@
Recently the SPICE/HeRALD collaboration provided a proof-of-principle for a cryogenic photomultiplier readout of a liquid helium volume [6]. The collaboration also kindly shared their published data with us in digital form which we use for our design. Our basic design is shown in Fig. 2. We have envisioned a hermetic copper vessel with a large gas space above a small cube (222 cm) where liquid He will collect when the vessel is filled with 6.63 liters (22.9 atm) of He at room temperature and lowered to a temperature of 1 K. Liquification will be achieved by attaching the vessel to an off-the-shelf, commercially obtainable closed-cycle 1 K cryocooler [7]. The cubical portion of the vessel will have a small quartz window through coated with tetraphenyl butadiene (TPB) so the phototube can measure the helium scintillation [8].
@@ 2 Expected Spectrum
My research group has estimated the measured neutron spectrum for a 1 yr SNOLAB exposure of the 222 cm liquid He volume [2]. We used the SNOLAB spectrum of Fig. 1 as an input spectrum and obtained the deposited-energy spectrum shown in Fig. 3
Our liquid He neutron detector design is appropriate for low neutron flux environments because of its high efficiency relative to conventional gaseous He tubes like those used by Beimer [9]. This is because of the much higher density of He in the liquid state at 1 K–0.0792 g/cm–compared to a 6 atm gas with density 7.410 g/cm at room temperature [10]. In fact, our 222 cm detector is expected register 3 times more neutron events above around 10 keV than the larger tubes (5 cm diameter; 15 cm length) of Beimer [9]. For typical gaseous tubes with fill pressures between 6–10 atm our design will be 64–107 times more efficient per volume at neutron energies larger than 10 keV. Below 10 keV the efficiency will quickly approach 100%.
The figure shows that for a 1 yr exposure we will produce an excellent measurement of the thermal neutron flux–given by the size of the peak near a deposited energy equal to the Q-value of the (n,p) reaction. It is also likely that a run period of that time will produce 3–4 counts in the high-energy neutron region above 2700 keV. Neutrons of those energies will constrain the overall fast flux present at the detector location. Similarly, the counts in the intermediate region, above around 266 keV neutron energy but below 1000 keV will create an important constraint in a region that has previously been difficult to measure.
@@ 3 Backgrounds to Neutron Measurement
When trying to measure low neutron fluxes with He–in a gaseous or liquid state–the dominant backgrounds are probably particles from the boundaries of the active volume. There are also lesser backgrounds from ambient gamma rays and radiogenic neutrons from the detector materials. Our liquid He design allows a large number of He atoms to be placed in a container with relatively low surface area compared to gaseous detectors. This generically limits the dominant background.
A good example is the Helium and Lead Observatory (HALO)–a gaseous He array operating at SNOLAB [11]. All together the array consists of around 0.75 m of He at 2.5 atm using special ultra-pure nickel tubes that reach a thorium contamination of 1 ppt [12]. That array will produce about 11,000 neutron counts above 1 keV a year, but around 140,000 background events. Our design with just 8 cm active volume will only produce 30 neutron events above 1 keV but could be expected to have around 115 background events in the same time period if we use highly radiopure copper (see below); a signal-to-noise ratio over two times better.
@@
Table 1 suggests to move to a scintillation readout that does not require a standard phototube or quartz window. This can be done using a bolometric Transition Edge Sensor (TES) or Microwave Kinetic Inductance Device (MKID) mounted on a silicon substrate inside the hermetic copper volume of our detector. Mounting just above the liquid He volume will allow good light collection without a quartz window. Silicon substrates can be made very clean and will not contribute much to the thorium contamination budget. With this modification we would expect to have a signal-to-noise ratio for neutrons above 1 keV of around 217 if we use ultra clean copper. This advancement would make for a high sensitivity neutron flux measurement even in very low background environments. Furthermore, the resolution and energy distributions of the signal and background can be taken into account for further improvements in the measurement.
4 Acknowledgements

### v2

# Liquid 3He Detectors for Neutrons

@@ Abstract
The 3He(n,p) process is excellent for neutron detection between thermal and 4 MeV because of the high cross section and near-complete energy transfer from the neutron to the proton. Traditional gaseous 3He detectors using this process typically have high levels of radiogenic backgrounds so that they cannot measure the small neutron fluxes present in underground labs for dark matter experiments. I propose a cryogenic liquid 3He detector that can be designed with tiny radiogenic backgrounds and efficiently measure neutron fluxes in low-flux environments.
@@
Low background underground environments, like SNOLAB, are home to many dark matter searches due to their minimization of non-signal backgrounds. The underground environments typically mitigate neutrons from cosmic rays (cosmogenic) very well, but there remains a neutron flux in the environment, primarily due to radiogenic sources–spontaneous fission or (,n) from the surrounding material [1]. The levels of these backgrounds are so low that they are difficult to measure, but still pose a threat to dark matter detectors as a background. My work aims to develop a prototype liquid 3He scintillation detector to measure the neutron backgrounds in low flux environments. The high cross section reaction 3He(n,p)3H in combination with the high 3He number density in the liquid state will allow these low fluxes to be measured by this prototype [2]. Measuring energy deposits as low as 500 eV in liquid 3He has been previously demonstrated [3] including work from the ULTIMA collaboration explicitly searching for low-energy recoils from dark matter [4]. Experiments have also specifically observed the 3He(n,p) process in liquid 3He [5, 6]. The scintillation of 3He should be similar to that of 4He and scintillation yield has been measured for liquid 4He [7] and reported in [2]. In the neutron energy range below 5 MeV the product proton travels less than 0.1 mm and thus very nearly all of the stopping energy of the proton and triton is deposited in the liquid volume.
@@
Figure 1 shows the neutron flux as a function of energy at SNOLAB. It was estimated from a combination of studies done in the past. First was a Monte Carlo approach targeting radiogenic neutrons from the cavern walls [8], this constrained the shape of the distribution for 6–10 MeV neutrons. The overall normalization for the “fast” neutrons was taken from the SNOLAB handbook [9]–the value was 4000 n/m2/day. The thermal flux–measured for SNO–is taken to come from a Maxwell-Boltzman spectrum with overall normalization that comes from measurement. In this case the measurement is that of Browne [10] and is quoted with the mean value first followed by the statistical and systematic uncertainties respectively: 4144.949.8105.3 n/m2/day. The intervals on each of the types of uncertainty were assumed Gaussian and symmetric. The flux in the area between the thermal and high-energy measurements is generally unmeasured and has been interpolated with a simple power-law function in Fig. 1, . Our prototype detector will measure the flux in this energy region which is especially important for low-mass dark matter searches.
@@
Recently the SPICE/HeRALD collaboration provided a proof-of-principle for a cryogenic photomultiplier readout of a liquid helium volume [7]. The collaboration also kindly shared their published data with us in digital form which we use for our design. Other collaborations like QUEST-DMC, and MACHe3 have also proposed the use of liquid 3He directly for dark matter searches [11, 12, 13]. Our basic design is shown in Fig. 2. We have envisioned a hermetic copper vessel with a large gas space above a small cube (222 cm3) where liquid 3He will collect when the vessel is filled with 6.63 liters (22.9 atm) of 3He at room temperature and lowered to a temperature of 1 K. Liquification will be achieved by attaching the vessel to an off-the-shelf, commercially obtainable closed-cycle 1 K cryocooler [14]. The cubical portion of the vessel will have a small quartz window through coated with tetraphenyl butadiene (TPB) so the phototube can measure the helium scintillation [15].
@@ 2 Expected Spectrum
My research group has estimated the measured neutron spectrum for a 1 yr SNOLAB exposure of the 222 cm3 liquid 3He volume [2]. We used the SNOLAB spectrum of Fig. 1 as an input spectrum and obtained the deposited-energy spectrum shown in Fig. 3
Our liquid 3He neutron detector design is appropriate for low neutron flux environments because of its high efficiency relative to conventional gaseous 3He tubes like those used by Beimer [16]. This is because of the much higher density of 3He in the liquid state at 1 K–0.0792 g/cm3–compared to a 6 atm gas with density 7.410-4 g/cm3 at room temperature [17]. In fact, our 222 cm3 detector is expected register 3 times more neutron events above around 10 keV than the larger tubes (5 cm diameter; 15 cm length) of Beimer [16]. For typical gaseous tubes with fill pressures between 6–10 atm our design will be 64–107 times more efficient per volume at neutron energies larger than 10 keV. Below 10 keV the efficiency will quickly approach 100%.
Figure 3 shows that for a 1 yr exposure we will produce an excellent measurement of the thermal neutron flux–given by the size of the peak near a deposited energy equal to the Q-value of the (n,p) reaction. It is also likely that a run period of that time will produce 3–4 counts in the high-energy neutron region above 2700 keV. Neutrons of those energies will constrain the overall fast flux present at the detector location. Similarly, the counts in the intermediate region, above around 266 keV neutron energy but below 1000 keV will create an important constraint in a region that has previously been difficult to measure.
@@ 3 Backgrounds to Neutron Measurement
When trying to measure low neutron fluxes with 3He–in a gaseous or liquid state–the dominant backgrounds are probably particles from the boundaries of the active volume. There are also lesser backgrounds from ambient gamma rays and radiogenic neutrons from the detector materials. Our liquid 3He design allows a large number of 3He atoms to be placed in a container with relatively low surface area compared to gaseous detectors. This generically limits the dominant background.
A good example is the Helium and Lead Observatory (HALO)–a gaseous 3He array operating at SNOLAB [18]. All together the array consists of around 0.75 m3 of 3He at 2.5 atm using special ultra-pure nickel tubes that reach a thorium contamination of 1 ppt [19]. That array will produce about 11,000 neutron counts above 1 keV a year, but around 140,000 background events. Our design with just 8 cm3 active volume will only produce 30 neutron events above 1 keV but could be expected to have around 115 background events in the same time period if we use highly radiopure copper (see below); a signal-to-noise ratio over two times better.
@@
| other materials not directly contacting sensitive volume555measurements take from the QUEST-DMC project [11] | |||
| stainless steel | 959 | 4421 | - |
| aluminum | 8.76104 | 4.04105 | - |
| stycast | 3.15103 | 1.45104 | - |
@@
Table 1 suggests to move to a scintillation readout that does not require a standard phototube or quartz window. This can be done using a bolometric Transition Edge Sensor (TES) or Microwave Kinetic Inductance Device (MKID) mounted on a silicon substrate inside the hermetic copper volume of our detector. Mounting just above the liquid 3He volume will allow good light collection without a quartz window. Silicon substrates can be made very clean and will not contribute much to the thorium contamination budget. With this modification we would expect to have a signal-to-noise ratio for neutrons above 1 keV of around 217 if we use ultra clean copper. This advancement would make for a high sensitivity neutron flux measurement even in very low background environments. Furthermore, the resolution and energy distributions of the signal and background can be taken into account for further improvements in the measurement.
There are detectors–like those proposed by the QUEST-DMC collaboration–that may realize the ability to use superfluid 3He and implement some sort of heat readout [11, 12]. This readout in addition to the scintillation readout we propose might allow for particle discrimination–the ratio of scintillation to heat is typically different for heavier recoils. In that case significant background mitigation might be achievable in our measurement because we could select proton recoils and exclude electron or alpha recoils. In fact Winkelmann and collaborators [22] have shown that for proton and tritium recoils resulting from the 3He(n,p) reaction on superfluid 3He will produce less than 15% scintillation fraction whereas cosmic-ray muons and low-energy electrons will have around 21% and 23% scintillation fractions respectively. The possibility of heat measurement would increase the cost of the prototype explored here but may warrant further study.
4 Acknowledgments